determine inverse Laplace transform of a-c

10s (a) (s + 1)(s + 2)(s + 3) 2s+ 4s + 1 (b) (s + 1)(s + 2) s +1 (c) (s + 2)(s? + 2s + 5)
Answer #1

los (5+1) (5+2) (5+3) A + + (5+3) (st) (5+2) 10X(-1) for a put 5:1 (-1+2) (-1+3) - 10 fuck 8 put 5= -2 10x (-2) - (-2+1) (-2725² +45t A B (b). (5+1) + (5+2) (579)2 1572)3 (5+1) (Sta} 2(-1)2 4.(-1)+1 for a put sa-1 (172)3 2-4t A - + + (st) (sta) (5+2)+1 3 3 (SH) (572) (stoje (5+2) + 3 * 1 (5+1) (542) (-1) x2! 2 (St2) 2+1 (5+2) it + 7 2 + 3 I (572) 2! (5+) (579)1+ ((st2)2+1For put s=-2 (-2+) +193) (-27-32). f 1723) (-1791) (23) 2_ coe e 1 for 8 put 5 = -1-2i 1-1-2i+1) (-1-25 +1-21) Ya 12 5= four